# Optimal estimation for regression discontinuity design

### v1

@@ 2.1 Setting
Shape restrictions on the second derivatives are common in studies on honest inference in RD designs (e.g., Kolesár and Rothe, 2018; Imbens and Wager, 2019; Noack and Rothe, 2024). The restriction of bounded second derivatives aligns with local linear estimators, for example. Nevertheless, we focus on the Lipschitz class for two reasons. First, restrictions on the second derivatives are less transparent and more challenging to evaluate than the Lipschitz constraints, which bound the partial effects of the running variable on the outcome. Second, the bounded second derivative implies the bounded first derivative when the regression function is bounded. To see this, suppose the domain of $f$ is $\mathbb{R}$ and the absolute value of the second derivative $f^{\prime\prime}(x)$ is bounded by $C>0$, so that $f^{\prime}(x+u)>f^{\prime}(x)-Cu$ for $u>0$. Then, we obtain $f(x+\delta)-f(x)=\int_{0}^{\delta}f^{\prime}(x+u)du\geq f^{\prime}(x)\delta-C\delta^{2}/2$ for any $\delta>0$. If the range of $f$ is $[0,1]$, $f(x+\delta)-f(x)$ must be less than or equal to $1$. Consequently, the first derivative satisfies $f^{\prime}(x)\leq\delta^{-1}+C\delta/2$ for any $\delta>0$, which implies that $f^{\prime}(x)\leq\min_{\delta>0}(\delta^{-1}+C\delta/2)=\sqrt{2C}$. In other words, the absolute value of the first derivative is bounded by $\sqrt{2C}$ when the absolute value of the second derivative $f^{\prime\prime}(x)$ is bounded by $C$ and the range of $f$ is $[0,1]$. In this manner, the second derivative restriction is closely related to the Lipschitz constraint for bounded outcomes.
@@
The solution of (2.3) is also the minimax linear shrinkage estimator for bounded outcomes. Consider the estimation of $p_{0}$ under the assumption that $P(0\leq Y_{i}\leq 1)=1$ and $\bm{p}\in\mathcal{P}$, where $p_{i}=E[Y_{i}]$. We impose no additional assumptions on $Y_{i}$. Then the variance of $Y_{i}$ must be less than or equal to $p_{i}(1-p_{i})$ because we have
@@
where the inequality follows from $P(Y_{i}^{2}\leq Y_{i})=1$. Since the bias of a linear estimator is the same for bounded and binary outcomes, the worst-case MSE for bounded outcomes is equal to the worst-case MSE for binary outcomes. Hence, the solution of (2.3) is also the minimax linear shrinkage estimator when $Y_{i}\in[0,1]$ and $\bm{p}\in\mathcal{P}$.
@@ 2.2 Computing the worst-case MSE of a linear shrinkage estimator
Our goal is to obtain the weight vector $\bm{w}$ that minimizes the maximal MSE. First, we consider the maximization part of (2.4) for a given weight vector $\bm{w}\in\mathcal{W}$. We show that the maximization problem with the ($n+1$)-dimensional parameter $\bm{\theta}=(\theta_{0},\ldots,\theta_{n})^{\prime}$ can be simplified into a maximization problem with a single parameter $\theta_{0}$.
@@
In view of Lemma 2.1, we may consider the maximization of the MSE over $\bm{\theta}\in\Theta$ satisfying the following restriction
@@
By calculating the derivatives of the MSE, we can show that $\text{MSE}(\bm{w},\bm{\theta})$ is nondecreasing in $\theta_{j}$ under (2.5). To see this, observe that
@@
Because we have $\sum_{i\neq j}w_{i}\theta_{i}-\theta_{0}\geq\left(\sum_{i\neq j}w_{i}-1\right)\theta_{0}\geq 0$ for all $\bm{w}\in\mathcal{W}$ under (2.5), it follows from (2.6) that $\text{MSE}(\bm{w},\bm{\theta})$ is nondecreasing in $\theta_{j}$ under (2.5). This monotonicity of the MSE implies that $\text{MSE}(\bm{w},(\theta_{0},\theta_{1},\ldots,\theta_{n})^{\prime})$ is maximized by setting $\theta_{1},\ldots,\theta_{n}$ to their largest possible values satisfying the Lipschitz constraint for each fixed value of $\theta_{0}$.
@@
$\tilde{\bm{\theta}}(t)\equiv\left(\tilde{\theta}_{0}(t),\tilde{\theta}_{1}(t),\ldots,\tilde{\theta}_{n}(t)\right)^{\prime}\ \text{and}\ \tilde{\theta}_{i}(t)\equiv\min\{t+C\|R_{i}\|,1/2\}\ \text{for $i=0,1,\ldots,n$}$
@@
while $\tilde{\bm{\theta}}(\theta_{0})$ satisfies (2.5). We also have $\tilde{\bm{\theta}}(t)\in\Theta$ for any $t\in[-1/2,1/2]$ because $\tilde{\bm{\theta}}(t)$ satisfies $\tilde{\bm{\theta}}(t)\in[-1/2,1/2]^{n+1}$ and
$\left|\tilde{\theta}_{i}(t)-\tilde{\theta}_{j}(t)\right|\ \leq\ C\left|\|R_{i}\|-\|R_{j}\|\right|\ \leq\ C\|R_{i}-R_{j}\|.$
Hence, we can reduce the ($n+1$)-dimensional maximization problem in (2.4) to a one-dimensional problem with the single parameter $\theta_{0}$ as in the following theorem:
@@ 2.3 The minimax linear shrinkage estimator
Next, we derive the weight vector that minimizes the maximum MSE. The following two lemmas show that the optimal weight vector is nonincreasing and that the $i$-th element of the optimal weight vector is zero if $R_{i}$ is sufficiently far away from $R_{0}$.
@@
Lemma 2.2 shows that the optimal weight vector must be nonincreasing. In the proof of Lemma 2.2, we show that if $\bm{w}\in\mathcal{W}$ satisfies $w_{j}<w_{j+1}$, the maximum MSE can be reduced by swapping the positions of $w_{j}$ and $w_{j+1}$. By repeating this procedure until the weight vector becomes monotone, we can obtain $\tilde{\bm{w}}\in\mathcal{W}_{0}$ such that $\max_{\bm{\theta}\in\Theta}\text{MSE}(\tilde{\bm{w}},\bm{\theta})\leq\max_{\bm{\theta}\in\Theta}\text{MSE}(\bm{w},\bm{\theta})$. Lemma 2.3 shows that the $i$-th element of the optimal weight vector is zero if $C\|R_{i}\|\geq 1/2$. By calculating the derivative of $\text{MSE}(\bm{w},\tilde{\bm{\theta}}(\theta_{0}))$ with respect to $w_{i}$, we can show that $\text{MSE}(\bm{w},\tilde{\bm{\theta}}(\theta_{0}))$ is nondecreasing in $w_{i}$ when $C\|R_{i}\|\geq 1/2$ and hence, setting $w_{i}=0$ is optimal.
These two lemmas allow us to restrict our search space for the optimal $\bm{w}$ to nonincreasing vectors that place no weight on the observations with $C\|R_{i}\|\geq 1/2$. For notational simplicity, we assume without loss of generality that our sample includes observations with $C\|R_{i}\|<1/2$ only, so that $\mathcal{W}_{0}=\mathcal{W}_{1}$. Theorem 2.1 and Lemma 2.2 then imply that the minimax problem is reduced to
@@
We now present how one can numerically solve the minimax problem (2.8). We define $g(\bm{w};\theta_{0})\equiv\text{MSE}(\bm{w},\tilde{\bm{\theta}}(\theta_{0}))$ and $\overline{g}(\bm{w})\equiv\max_{\theta_{0}\in[-1/2,0]}g(\bm{w};\theta_{0})$. Because both $\bm{w}\mapsto\left(\sum_{i=1}^{n}w_{i}\theta_{i}-\theta_{0}\right)^{2}$ and $\bm{w}\mapsto\sum_{i=1}^{n}w_{i}^{2}\left(\frac{1}{4}-\theta_{i}^{2}\right)$ are convex for any $\bm{\theta}\in\Theta$, $g(\bm{w};\theta_{0})$ is also convex with respect to $\bm{w}$ for any $\theta_{0}\in[-1/2,0]$. Because the maximum of convex functions is also convex, $\overline{g}(\bm{w})$ is a convex function. Therefore, the minimax problem (2.8) becomes the following convex optimization problem with linear constraints:
@@
Hence, we may compute the optimal $\bm{w}$ by solving a linearly constrained convex optimization problem where its objective function can be evaluated by a scalar-valued grid search for the optimizing $\theta_{0}$.
@@
In the implementation in simulations and applications, we use a nonlinear optimization via augmented Lagrange method (Ghalanos and Theussl, 2015; Ye, 1987). Nevertheless, $g(\bm{w};\theta_{0})$ is a quadratic function in $\theta_{0}$ and $\overline{g}(\bm{w})$ has a closed-form expression. Let $u(\bm{w})\equiv\sum_{i=1}^{n}w_{i}$ and $k(\bm{w})\equiv\sum_{i=1}^{n}w_{i}\|R_{i}\|$. Then, $g(\bm{w};\theta_{0})$ can be written as
@@
Combining the two cases, $g(\bm{w};\theta_{0})$ is maximized at $\theta_{0}=-1/2$ if and only if the following inequality holds:
@@
In this remark, we return to the original setup introduced in Section 2.1, where we observe both the treated sample $\{Y_{i,+},R_{i,+}\}_{i=1}^{n_{+}}$ and the untreated sample $\{Y_{i,-},R_{i,-}\}_{i=1}^{n_{-}}$. We consider the estimation of $f(1,R_{0})-f(0,R_{0})$, which can be interpreted as the conditional average treatment effect (ATE) at the cutoff $R_{0}$. We may estimate the ATE by separetely constructing the aforementioned minimax linear shrinkage estimators for $f(1,R_{0})$ and $f(0,R_{0})$ using the treated and untreated samples respectively. Specifically, let $\hat{\bm{w}}_{+}$ and $\hat{\bm{w}}_{-}$ be the optimal weights that minimize the maximum MSEs among linear shrinkage estimators of $f(1,R_{0})$ and $f(0,R_{0})$. Then, we can estimate the conditional ATE $f(1,R_{0})-f(0,R_{0})$ using the following estimator:
@@
Note that this estimator does not minimize the maximum MSE for the ATE estimation among estimators that take the difference between two linear shrinkage estimators; the MSE for $f(1,R_{0})-f(0,R_{0})$ is not equal to the sum of the MSEs for $f(1,R_{0})$ and $f(0,R_{0})$. Nevertheless, Appendix B shows that we can still obtain results similar to Theorem 2.1 and Lemmas 2.2 and 2.3 at the cost of an additional grid search and a possible instability in the estimate. Specifically, the maximum MSE for the ATE can be calculated by simultaneously optimizing two parameters, $f(1,R_{0})$ and $f(0,R_{0})$.
@@ 3 Comparison with Gaussian-motivated estimators
Many existing studies consider minimax estimation problems for unbounded outcomes with known variance, primarily motivated by the Gaussian model. We compare our proposed estimator with a Gaussian-motivated minimax linear estimator when the underlying data generating process is the binary outcome model.
Following the existing minimax analysis in RD designs (Armstrong and Kolesár, 2018; Imbens and Wager, 2019), we consider the Gaussian-motivated estimator as the minimax estimator for an unbounded space of mean vectors with known variances under the Lipschitz constraint as in Section 2. Note that if the outcome $Y_{i}$ is normally distributed, that is, $Y_{i}\sim N(p_{i},\sigma_{i}^{2})$, the MSE of a linear estimator $\hat{p}_{0}(\bm{w})=\frac{1}{2}+\sum_{i=1}^{n}w_{i}\left(Y_{i}-\frac{1}{2}\right)$ with $\bm{w}\in\mathbb{R}^{n}$ is given by
@@
As a smoothness restriction, we impose the Lipschitz constraint where the parameter space is given by
@@
The minimax linear estimator is the solution of the following problem:
@@
We refer to the linear estimator that solves (3.1) as the Gaussian estimator. [^4] Note that this estimator is a minimax linear estimator without normality of $Y_{i}$ as long as variance is known and the parameter space is $\Theta_{g}$. Normality of $Y_{i}$ is exploited for finite-sample valid inference based on a linear estimator. This above minimax problem (3.1) differs from the original binary-outcome problem (2.4) in three aspects. First, the minimum in (3.1) is considered among all linear estimators, including those with negative weights. Second, the parameter space in (3.1) is unbounded. Lastly, but most importantly, the variance in (3.1) does not depend on the parameter $\bm{\theta}$, and hence the maximum MSE is attained at the parameter values that maximize the squared bias.
In Appendix D, we derive the form of the optimal weights that solve the minimax problem (3.1) by an application of the results in Donoho (1994) to our Gaussian setting. We show that the optimal weights satisfy $\sum_{i=1}^{n}w_{i}=1$ and $w_{i}\geq 0$ for all $i$. Hence, the minimax problem (3.1) can be solved by minimizing the maximum MSE on $\mathcal{W}$. More specifically, the Gaussian estimator is obtained by solving the following quadratic program:

### v2

@@ 2.1 Setting
Shape restrictions on second derivatives are common in studies of honest inference in RD designs (e.g., kolesar2018discrete; Imbens.Wager2019; noack_bias_aware_2024). One example is imposing bounds on second derivatives, which aligns with local linear estimators. We focus on the Lipschitz class for two reasons. First, restrictions on the second derivatives are less transparent and more challenging to evaluate than the Lipschitz constraints, which bound the partial effects of the running variable on the outcome. Second, the bounded second derivative implies the bounded first derivative when the regression function is bounded. To see this, suppose that the domain of $f$ is $\mathbb{R}$ and the absolute value of the second derivative $f^{\prime\prime}(x)$ is bounded by $C>0$, so that $f^{\prime}(x+u)\geq f^{\prime}(x)-Cu$ for $u>0$. Then, we obtain $f(x+\delta)-f(x)=\int_{0}^{\delta}f^{\prime}(x+u)du\geq f^{\prime}(x)\delta-C\delta^{2}/2$ for any $\delta>0$. If the range of $f$ is $[0,1]$, $f(x+\delta)-f(x)$ must be less than or equal to $1$. Consequently, the first derivative satisfies $f^{\prime}(x)\leq\delta^{-1}+C\delta/2$ for any $\delta>0$, which implies that $f^{\prime}(x)\leq\min_{\delta>0}(\delta^{-1}+C\delta/2)=\sqrt{2C}$. Similarly, we have $f^{\prime}(x)\geq-\sqrt{2C}$. In other words, the absolute value of the first derivative is bounded by $\sqrt{2C}$ when the absolute value of the second derivative $f^{\prime\prime}(x)$ is bounded by $C$ and the range of $f$ is $[0,1]$. Thus, the second-derivative restriction is closely related to the Lipschitz constraint for bounded outcomes.
@@
The solution to (2.3) is also a minimax linear shrinkage estimator for bounded outcomes. Consider the estimation of $p_{0}$ under the assumption that $P(0\leq Y_{i}\leq 1)=1$ and $\bm{p}\in\mathcal{P}$, where $p_{i}=E[Y_{i}]$. We impose no additional assumptions on $Y_{i}$. Then, the variance of $Y_{i}$ must be less than or equal to $p_{i}(1-p_{i})$ because we have
@@
where the inequality follows from $P(Y_{i}^{2}\leq Y_{i})=1$. Since the bias of a linear estimator is the same for bounded and binary outcomes, the worst-case MSE for bounded outcomes is equal to the worst-case MSE for binary outcomes. Hence, the solution to (2.3) is also a minimax linear shrinkage estimator when $Y_{i}\in[0,1]$ and $\bm{p}\in\mathcal{P}$.
Remark 2.4.
Our theoretical analysis extends to a class of Hölder continuous functions:
[…]
@@ 2.2 Computing the worst-case MSE of a linear shrinkage estimator
Our goal is to obtain a weight vector $\bm{w}$ that minimizes the maximum MSE. First, we consider the maximization part of (2.4) for a given weight vector $\bm{w}\in\mathcal{W}$. Note that the objective function $\text{MSE}(\bm{w},\bm{\theta})$ is generally nonconcave in $\bm{\theta}=(\theta_{0},\ldots,\theta_{n})^{\prime}$, as the squared bias $\left(\sum_{i=1}^{n}w_{i}\theta_{i}-\theta_{0}\right)^{2}$ is convex in $\bm{\theta}$, whereas the variance $\sum_{i=1}^{n}w_{i}^{2}\left(\frac{1}{4}-\theta_{i}^{2}\right)$ is concave in $\bm{\theta}$. Nevertheless, we show that this nonconvex ($n+1$)-dimensional optimization problem can be reduced to an optimization problem over a single parameter $\theta_{0}$, and is therefore computationally tractable.
@@
In view of Lemma 2.1, we may consider the maximization of the MSE over $\bm{\theta}\in\Theta$ subject to the following restriction:
@@
By calculating the derivatives of the MSE, we can show that $\text{MSE}(\bm{w},\bm{\theta})$ is nondecreasing in $\theta_{j}$ under (2.6). To see this, note that
@@
Because we have $\sum_{i\neq j}w_{i}\theta_{i}-\theta_{0}\geq\left(\sum_{i\neq j}w_{i}-1\right)\theta_{0}\geq 0$ for all $\bm{w}\in\mathcal{W}$ under (2.6), it follows from (2.7) that $\text{MSE}(\bm{w},\bm{\theta})$ is nondecreasing in $\theta_{j}$ under (2.6). This monotonicity of the MSE implies that $\text{MSE}(\bm{w},(\theta_{0},\theta_{1},\ldots,\theta_{n})^{\prime})$ is maximized by setting $\theta_{1},\ldots,\theta_{n}$ to the largest possible values that satisfy the Lipschitz constraint for each fixed value of $\theta_{0}$.
@@
$\tilde{\bm{\theta}}(t)\equiv\left(\tilde{\theta}_{0}(t),\tilde{\theta}_{1}(t),\ldots,\tilde{\theta}_{n}(t)\right)^{\prime}\ \text{and}\ \tilde{\theta}_{i}(t)\equiv\min\{t+C\|R_{i}\|,1/2\}\ \text{for $i=0,1,\ldots,n$,}$
@@
Also, $\tilde{\bm{\theta}}(t)\in\Theta$ for any $t\in[-1/2,1/2]$ because $\tilde{\bm{\theta}}(t)$ satisfies $\tilde{\bm{\theta}}(t)\in[-1/2,1/2]^{n+1}$ and
$\left|\tilde{\theta}_{i}(t)-\tilde{\theta}_{j}(t)\right|\ \leq\ C\left|\|R_{i}\|-\|R_{j}\|\right|\ \leq\ C\|R_{i}-R_{j}\|,$
where the second inequality follows from the reverse triangle inequality. Hence, we can reduce the ($n+1$)-dimensional maximization problem in (2.4) to a one-dimensional problem with a single parameter $\theta_{0}$, as in the following theorem.
@@ 2.3 The minimax linear shrinkage estimator
Next, we derive a weight vector that minimizes the maximum MSE. The following two lemmas show that the optimal weight vector is nonincreasing and that the $i$th element of the optimal weight vector is zero if $R_{i}$ is sufficiently far away from $R_{0}$.
@@
Lemma 2.2 shows that the optimal weight vector must be nonincreasing. In the proof of Lemma 2.2, we show that if $\bm{w}\in\mathcal{W}$ satisfies $w_{j}<w_{j+1}$, then the maximum MSE can be reduced by swapping the positions of $w_{j}$ and $w_{j+1}$. By repeating this procedure until the weight vector becomes monotone, we can obtain $\tilde{\bm{w}}\in\mathcal{W}_{0}$ such that $\max_{\bm{\theta}\in\Theta}\text{MSE}(\tilde{\bm{w}},\bm{\theta})\leq\max_{\bm{\theta}\in\Theta}\text{MSE}(\bm{w},\bm{\theta})$. Lemma 2.3 shows that the $i$th element of the optimal weight vector is zero if $C\|R_{i}\|\geq 1/2$. By calculating the derivative of $\text{MSE}(\bm{w},\tilde{\bm{\theta}}(\theta_{0}))$ with respect to $w_{i}$, we can show that $\text{MSE}(\bm{w},\tilde{\bm{\theta}}(\theta_{0}))$ is nondecreasing in $w_{i}$ when $C\|R_{i}\|\geq 1/2$, and hence, setting $w_{i}=0$ is optimal.
These two lemmas allow us to restrict our search space for the optimal $\bm{w}$ to nonincreasing vectors that place no weight on the observations with $C\|R_{i}\|\geq 1/2$. For notational simplicity, without loss of generality, we assume that our sample includes observations with $C\|R_{i}\|<1/2$ only, so that $\mathcal{W}_{0}=\mathcal{W}_{1}$. Theorem 2.1 and Lemma 2.2 then imply that the minimax problem is reduced to
@@
We now present a method for numerically solving the minimax problem (2.9). We define $g(\bm{w};\theta_{0})\equiv\text{MSE}(\bm{w},\tilde{\bm{\theta}}(\theta_{0}))$ and $\overline{g}(\bm{w})\equiv\max_{\theta_{0}\in[-1/2,0]}g(\bm{w};\theta_{0})$. Because both $\bm{w}\mapsto\left(\sum_{i=1}^{n}w_{i}\theta_{i}-\theta_{0}\right)^{2}$ and $\bm{w}\mapsto\sum_{i=1}^{n}w_{i}^{2}\left(\frac{1}{4}-\theta_{i}^{2}\right)$ are convex for any $\bm{\theta}\in\Theta$, $g(\bm{w};\theta_{0})$ is also convex with respect to $\bm{w}$ for any $\theta_{0}\in[-1/2,0]$. As the maximum of convex functions is also convex, $\overline{g}(\bm{w})$ is a convex function. Therefore, the minimax problem (2.9) becomes the following convex optimization problem with linear constraints:
@@
We use nonlinear optimization via the augmented Lagrange method (package_rsolnp; Ye_1987) in the implementation in simulations and applications.
@@
In the implementation, we compute $\bar{g}(\bm{w})$ by conducting a scalar-valued grid search to optimize $\theta_{0}$. Nevertheless, $g(\bm{w};\theta_{0})$ is a quadratic function in $\theta_{0}$ and $\overline{g}(\bm{w})$ has a closed-form expression. Let $u(\bm{w})\equiv\sum_{i=1}^{n}w_{i}$ and $k(\bm{w})\equiv\sum_{i=1}^{n}w_{i}\|R_{i}\|$. Then, $g(\bm{w};\theta_{0})$ can be written as
@@
Combining these two cases, $g(\bm{w};\theta_{0})$ is maximized at $\theta_{0}=-1/2$ if and only if the following inequality holds:
@@
In this remark, we return to the original setup introduced in Section 2.1, where we observe both the treated sample $\{Y_{i,+},R_{i,+}\}_{i=1}^{n_{+}}$ and the untreated sample $\{Y_{i,-},R_{i,-}\}_{i=1}^{n_{-}}$. We consider the estimation of $f(1,R_{0})-f(0,R_{0})$, which can be interpreted as the conditional average treatment effect (ATE) at the cutoff $R_{0}$. We can estimate the ATE by separately constructing the minimax linear shrinkage estimators for $f(1,R_{0})$ and $f(0,R_{0})$ using the treated and untreated samples, respectively. Specifically, let $\hat{\bm{w}}_{+}$ and $\hat{\bm{w}}_{-}$ be the optimal weights minimizing the maximum MSEs among the linear shrinkage estimators for $f(1,R_{0})$ and $f(0,R_{0})$. We can then estimate the conditional ATE $f(1,R_{0})-f(0,R_{0})$ using the following estimator:
@@
Note that this estimator does not minimize the maximum MSE for the ATE estimation among the estimators that take the difference between two linear shrinkage estimators; the MSE for $f(1,R_{0})-f(0,R_{0})$ is not equal to the sum of the MSEs for $f(1,R_{0})$ and $f(0,R_{0})$. In Appendix B, we consider the joint optimization of the weight vectors $\bm{w}$$_{+}$ and $\bm{w}$$_{-}$ and obtain results similar to those of Theorem 2.1 and Lemmas 2.2 and 2.3. Specifically, the maximum MSE for the ATE can be calculated by simultaneously optimizing two parameters, $f(1,R_{0})$ and $f(0,R_{0})$. Moreover, the optimal weight vectors are nonincreasing in the distance from the cutoff $R_{0}$. Although joint optimization of $\bm{w}$$_{+}$ and $\bm{w}$$_{-}$ is thus possible, it may require a two-dimensional grid search to calculate the worst-case MSE and potentially introduce instability in the resulting ATE estimates. Therefore, we use the separately optimized weights $\hat{\bm{w}}_{+}$ and $\hat{\bm{w}}_{-}$ to estimate the ATE in our simulations and empirical application.
@@ 3 Comparison with Gaussian-motivated estimators
Many existing studies have considered minimax estimation problems for unbounded outcomes with known variances, primarily motivated by the Gaussian model. We compare our proposed estimator with a Gaussian-motivated minimax linear estimator when the underlying data-generating process is a binary-outcome model.
As a Gaussian-motivated estimator, we consider the minimax linear estimator for an unbounded space of mean vectors with known variances under a smoothness restriction, following the existing minimax analysis in RD designs (Armstrong.Kolesar2018; Imbens.Wager2019). Note that if the outcome $Y_{i}$ is normally distributed, that is, $Y_{i}\sim N(p_{i},\sigma_{i}^{2})$, the MSE of a linear estimator $\hat{p}_{0}(\bm{w})=\frac{1}{2}+\sum_{i=1}^{n}w_{i}\left(Y_{i}-\frac{1}{2}\right)$ with $\bm{w}\in\mathbb{R}^{n}$ is given by
@@
As a smoothness restriction, we impose the Lipschitz constraint, leading to the following parameter space:
@@
The minimax linear estimator is a solution of the following problem:
@@
We refer to the linear estimator that solves (3.1) as the Gaussian estimator. [^5] Note that this estimator is a minimax linear estimator without normality of $Y_{i}$ as long as the variance is known and the parameter space is $\Theta_{g}$. The normality of $Y_{i}$ is exploited for finite-sample valid inference based on a linear estimator. The above minimax problem (3.1) differs from the original binary-outcome problem (2.4) in three respects. First, the minimum in (3.1) is considered among all linear estimators, including those with negative weights. Second, the parameter space in (3.1) is unbounded. Finally, but most importantly, the variance in (3.1) does not depend on the parameter $\bm{\theta}$; hence, the maximum MSE is attained at the parameter values that maximize the squared bias.
In Appendix D, we derive the form of the optimal weights that solve the minimax problem (3.1) by applying the results of donoho1994 to our Gaussian setting. We show that the optimal weights satisfy $\sum_{i=1}^{n}w_{i}=1$ and $w_{i}\geq 0$ for all $i$. Hence, the minimax problem (3.1) can be solved by minimizing the maximum MSE over $\mathcal{W}$. Specifically, the Gaussian estimator is obtained by solving the following quadratic program: